OpenStudy (anonymous):

(medal for the person who helps) Indicate in standard form the equation of the line passing through the given points. M(0, 6), N(6, 0)

OpenStudy (anonymous):

Standard form is Ax+By=C The first thing you want to do is find the slope: m=(y-y1)/(x-x1) Let's say M(0,6) is (x,y) and N(6,0) is (x1,y1) so m=(6-0)/(0-6)=-1 We can start with point slope form which is (y-y1)=m(x-x1) We can fill in either x,y or x1,y1 y-0=-1(x-6) y=-x+6 Then we want to move all the terms with x and y to one side and the constants(numbers by themselves) to the other side and you'll get your answer. x+y=6

OpenStudy (anonymous):

so this is the slope form ?